OpenStudy (anonymous):

4. What is the undefined value of 3b^2+13b+4/b+4? (1 point)

jimthompson5910 (jim_thompson5910):

You can't divide by zero. So the denominator can't be equal to zero. If it were, then b + 4 =0, solve for b b+4 = 0 b+4-4 = 0-4 b = -4 So if b = -4, then the denominator is zero, which makes the entire expression undefined.